In a laboratory, the count of bacteria in a certain experiment was increasing at the rate of 2.5% per hour. Find the bacteria at the end of 2 hours if the count was initially 5,06,000. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The initial count of bacteria is given as 5,06,000.
Bacteria at the end of 2 hours=506000(1 +2.5/100 )^2
= 506000(1 +1/40 )^2
= 506000(1 + 41/40 )^2
= 531616.25
∴ The count of bacteria at the end of 2 hours will be 5,31,616 (approx.)

​Thanks